OpenStudy (anonymous):

Solve the system x - y = 3 and 3x - 8y = 9. What is the value of y? 3 0 6 -2 @marissalovescats

OpenStudy (anonymous):

@satellite73 would I work this one the same way as the last one?

OpenStudy (anonymous):

\[x-y=3\\3x-8y=9\] if you want you can multiply the first equation by \(-3\) and get \[-3x+3y=-9\\ 3x-8y=9\]

OpenStudy (texaschic101):

you can use substitution or elimination

OpenStudy (marissalovescats):

Use elimination to get rid of the x to solve for y so times the whole 1st equation by -3 and tell me what you get

OpenStudy (anonymous):

you can't just add in this case because nothing will drop out of you do but if you multiply the first equation all the way across by \(-3\) then the \(-3x\) and the \(3x\) will add up to zero

OpenStudy (anonymous):

9? @marissalovescats

OpenStudy (texaschic101):

or sub 3 - y in for x in the 2nd equation

OpenStudy (anonymous):

\[-3x+3y=-9\\ 3x-8y=9\] now if you add them the \(x\)'s go away and you get \[-5y=0\]which tells you \(y=0\)

OpenStudy (marissalovescats):

Yup I got 0 ^^

OpenStudy (anonymous):

o some how I ended up with 3, but I see what I did I wrote it messy and used a different number thanks! both of you @satellite73 @marissalovescats